OpenStudy (anonymous):

On which of the following intervals is the function f(x) = 4 cos(2x − π) decreasing? x = pi over 2 to x = π x = 0 to x = pi over 2 x = pi over 2 to x = 3 pi over 2 x = π to x = 3 pi over 2

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (anonymous):

is it C?

OpenStudy (loser66):

How can you get it?

OpenStudy (anonymous):

graphing I guess?

OpenStudy (loser66):

ok, let see, you choose c, it means [pi/2, 3pi/2] now plug into the original one at x = pi/2, 4 cos (2x -pi ) = 4 cos ( 2(pi/2) -pi) = 4 cos 0 = 4 next is x = 3pi/2, 4 cos (2x - pi) = 4 cos (3pi -pi) = 4 cos (2pi) = 4. Now, a point between them: I pick pi 4cos (2x -pi) = 4 cos (2pi - pi) = 4 cos pi = -4 now you see. 

OpenStudy (loser66):

but at pi it is -4, hence it must be down from 4 to -4